OpenStudy (anonymous):

solve for x: -3|x-3| = -6

OpenStudy (anonymous):

x = 1, x = -1 x = 1, x = 5 x = 0, x = 5 no solutions

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@iPwnBunnies

OpenStudy (anonymous):

@IMStuck

OpenStudy (jdoe0001):

\(\bf -3|x-3| = -6\implies |x-3| =\cancel{ \cfrac{-6}{-3} }\implies |x-3| =2 \\ \quad \\ |x-3| =2\to \begin{cases} +(x-3)=2 \\ \quad \\ \bf -(x-3)=2 \end{cases}\)

OpenStudy (jdoe0001):

absolute value means, two scenarios, thus two values for "x", solve each to get the two "x values"